OpenStudy (anonymous):

how many days are there in 240,000 seconds?

OpenStudy (luigi0210):

Hello @kyia1 Welcome to openstudy :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Divide 240,000 by number of seconds in a day no of seconds in a day = 24*60*60 = 86400 no of days = 240,000 /86400=2.7 days approx

OpenStudy (anonymous):

how many inches are there in a football field (100yds) ?

OpenStudy (anonymous):

a small herd of cattle consumes fourteen bales of hay in two weeks . how many bales will this herd consume in a year?

OpenStudy (anonymous):

a=h(a+b)\2 ,solve for h